# Scientific Large Language Models: A Survey on Biological & Chemical Domains

### v1

@@ 3.3. Evaluation
Capabilities. The evaluation of LLMs is often structured around Bloom’s taxonomy (Krathwohl, 2002; Forehand, 2010), which outlines six cognitive levels of educational learning objectives: Remember, Understand, Apply, Analyze, Evaluate and Create. More recently, SciEval (Sun et al., 2023) has suggested assessing scientific LLMs across four dimensions: basic knowledge, knowledge application, scientific calculation, and research ability. Each dimension aligns with one or more cognitive levels in Bloom’s taxonomy. While these frameworks offer cognitive and research perspectives, our focus is to gauge the depth of knowledge grasped by LLMs, ideally comparable to human comprehension. Therefore, this survey adopts an approach (named KnowEval) that categorizes the capabilities of Text-Sci-LLMs based on the complexity of scientific knowledge, differentiating between Pre-college, College, and Post-college levels, as illustrated in Fig. 7.
@@
  • College Level. Scientific knowledge at the college level is typically more specialized, abstract, and necessitates logical reasoning and proof rather than mere memorization. It undergoes frequent updates aligned with scientific and industrial advancements, posing a challenge for LLMs to grasp and adapt to these continual changes. The master of specialized knowledge is designated as the college level of LLMs, which typically aligns with the Apply and Analyze levels in Bloom’s taxonomy, as well as the knowledge application and the scientific calculation in SciEval. Most evaluation benchmarks such as PubMedQA (Jin et al., 2019) and SciQ (Johannes Welbl, 2017) focus on this level.
@@
Figure 7. Comparsion of evaluation principle in (a) SciEval, (b) Bloom, and (c) KnowEval (this survey).
@@
When the ground-truth is absent, one could sort to experts or strong agents like GPT-4 (OpenAI, 2023) for assessing LLMs. Human evaluation plays a pivotal role in assessing the quality and accuracy of model-generated results, as it closely aligns with real-world application scenarios and provides comprehensive and precise feedback. However, due to extensive human involvement, this process is time-consuming and labor-intensive. Therefore, there is a growing tendency that employ GPT-4 as a “judge” to evaluate different models (Bubeck et al., 2023). Overall, these metrics are often used complementarily in various NLP tasks to provide a comprehensive evaluation of LLMs.
@@ 4. Molecular Large Language Models
(Sec. 4.1),text width=4em [Encoder-only, text width= 5.5em [BERT-based: SMILES-BERT (Wang et al., 2019)   MTL-BERT (Zhang et al., 2022b)  rxnfp-BERT (Schwaller et al., 2021)
Mol-BERT (Li and Jiang, 2021)  MolFormer (Ross et al., 2022)   MolRoPE-BERT (Liu et al., 2023d) ] [RoBERTa-based: ChemBERTa (Chithrananda et al., 2020)  MFBERT (Abdel-Aty and Gould, 2022)
SELFormer (Yüksel et al., 2023)   Semi-RoBERTa (Tran and Ekenna, 2023)  ] [Integrating 2D or 3D graphs: GROVER (Rong et al., 2020)  MAT (Maziarka et al., 2020)  MG-BERT (Zhang et al., 2021)
KPGT (Li et al., 2022)  AGBT (Chen et al., 2021)   Molformer (Wu et al., 2023)  Uni-Mol (Zhou et al., 2023a)  ] ] [Decoder-only, text width= 5.5em [GPT-based: MolGPT (Bagal et al., 2021)  SMILESGPT (Adilov, 2021)  cMolGPT (Wang et al., 2023f)   Taiga (Mazuz et al., 2023)  iupacGPT (Cho et al., 2023) ] ] [Encoder-decoder, text width=5.5em [Vanilla Transformer-based:
@@
(Sec. 4.2),text width=4em [Pre-training, text width=4em [ZINC (Sterling and Irwin, 2015)  PubChem (Wang et al., 2012)  USPTO (Schneider et al., 2016)  PCQM4M (Hu et al., 2021; Ying et al., 2021)  GEOM (Axelrod and Gomez-Bombarelli, 2022)
MolTLU (Beaini et al., 2023)  ChEMBL (Gaulton et al., 2012)   DrugBank (Wishart et al., 2018)  GDB-17 (Ruddigkeit et al., 2012)  ExCAPE-DB (Sun et al., 2017)  ] ] [Benchmarks, text width=4em [MoleculeNet (Wu et al., 2018)  MARCEL (Zhu et al., 2023)  GuacaMol (Brown et al., 2019)   MOSES (Polykovskiy et al., 2020)  ADMETlab 2.0  (Xiong et al., 2021) ] ] ] [Evaluation
[…]
@@ 4.1. Models
As Transformer (Vaswani et al., 2017) can effectively model sequential data, numerous Transformer-based Mol-LLMs have been proposed for predicting molecular properties and structures. Here we categorize Mol-LLMs into three distinct types based on their specific architectures: encoder-only, decoder-only, and encoder-decoder models, as shown in Table 3.
@@ 4.1.1. Encoder-only Models
Encoder-only models focus on understanding and interpreting the input molecules, making them well-suited for tasks that require a deep comprehension of molecular structures and properties. A typical example is SMILES-BERT (Wang et al., 2019) that leverages the BERT architecture to interpret SMILES representations of molecules. It undergoes pre-training on an extensive dataset of unlabeled data through masked SMILES recovery tasks, enabling it to learn nuanced patterns and relationships within molecular structures. Nevertheless, the fine-tuning phase for individual downstream tasks is inefficient. MTL-BERT (Zhang et al., 2022b) mitigates this limitation by employing a multitask learning framework, which substantially increases the accuracy of predictions for multiple molecular properties. The rxnfp-BERT (Schwaller et al., 2021) model draws on the principles of BERT to process and learn from chemical reactions. It is designed to generate reaction fingerprints, which are representations of chemical reactions used for various predictive tasks. However, these models predominantly concentrate on contextual information in molecular sequences while giving less attention to molecular substructures, such as functional groups. Mol-BERT (Li and Jiang, 2021) extracts atomic-level and substructural features centered on the current atom. MoLFormer (Ross et al., 2022) further enhances this approach by integrating relative position encoding and rotational position embedding, thereby facilitating the comprehension of spatial relationships between atoms in a molecule. MolRoPE-BERT (Liu et al., 2023d) augments this method by employing Rotary Position Embedding (RoPE) to efficiently encode positional information in SMILES, aiming to provide a more robust understanding of molecules.
@@
MTL-BERT (Zhang et al., 2022b) 2022.12 - BERT ChEMBL Prop. pred.
MolRoPE-BERT (Liu et al., 2023d) 2023.01 - BERT ZINC-15, ChEMBL 27 Prop. pred. $\times$
Molformer (Wu et al., 2023) 2023.01 - Trans. enc. ZINC, PubChem Prop. pred.
@@
cMolGPT (Wang et al., 2023f) 2023.05 - GPT MOSES Mol. gen.
@@
The above BERT-based models integrate a vast amount of unlabeled molecular data through a self-supervised learning strategy, but they are based on the sequential format of molecule SMILES or SELFIES, overlooking crucial graph structures and three-dimensional (3D) chemical information. The oversight of graph structures can be effectively addressed by integrating graphs with Transformers, such as GROVER (Rong et al., 2020), MAT (Maziarka et al., 2020), MG-BERT (Zhang et al., 2021), and KPGT (Li et al., 2022). For example, MG-BERT (Zhang et al., 2021) combines GNN’s local message passing with BERT’s architecture, enhancing molecular feature extraction and utilizing a unique pretraining strategy involving atom masking. Certainly, GNNs alone can deal with molecule graph structures, and representative works include KCL (Fang et al., 2022), KANO (Fang et al., 2023d), GS-Meta (Zhuang et al., 2023b), and iMoLD (Zhuang et al., 2023a). Moreover, several approaches have been employed to model 3D molecules, including AGBT (Chen et al., 2021), Molformer (Wu et al., 2023), and Uni-mol (Zhou et al., 2023a), and DeepSorption (Cui et al., 2023), showcasing the evolution and diversification of computational methods for molecular modeling. For example, DeepSorption leverages a SE(3)-equivalent Transformer to effectively model the 3D structures of molecular crystals, leading to accurate prediction of gas adsorption. Graph and 3D structure-based molecular representation learning is out the scope of this survey and we refer interested readers to the surveys (Guo et al., 2023; Xia et al., 2023).
@@ 4.1.2. Decoder-only Models
In recent years, decoder-only architectures like GPT have revolutionized molecular generation in cheminformatics. Mol-LLMs based on the GPT architecture, including MolGPT (Bagal et al., 2021), SMILESGPT (Adilov, 2021), iupaccGPT (Cho et al., 2023), cMolGPT (Wang et al., 2023f), and Taiga (Wang et al., 2022b), primarily use SMILES strings as input, effectively addressing the challenge of navigating the vast chemical space. These models are pivotal in drug discovery and material science, enabling the synthesis of molecules with specific properties.
MolGPT (Bagal et al., 2021), a pioneer in utilizing GPT for molecular generation, uses conditional training for property optimization. It stands out for its efficiency and effectiveness in molecular modeling and drug discovery, showing strong control over multiple properties for accurate generation. Moreover, both SMILESGPT and iupacGPT are based on the GPT-2 architecture. SMILESGPT (Adilov, 2021) is a causative Transformer tailored for drug discovery, training on SMILES notation with BPE, and employing special embedding layers for effective molecular representation learning. iupacGPT (Cho et al., 2023) diverges by using the more readable IUPAC nomenclature, excelling in learning hierarchical and inner-sequence relationships between atoms and chemical groups. The cTransformer (Wang et al., 2022b) and cMolGPT (Wang et al., 2023f) models are specifically developed for the purpose of generating molecules based on given conditions. cTransformer (Wang et al., 2022b) synergizes the Transformer architecture with conditional embeddings, allowing for target-specific molecule generation and fine-tuning with target-specific data. cMolGPT (Wang et al., 2023f) represents an advancement in this domain by generating target-specific compounds by utilizing target embeddings as keys and values within its multi-head attention, adeptly producing molecules active against designated targets. Taiga (Mazuz et al., 2023) exemplifies the combination of GPT and reinforcement learning (RL) methodologies. It employs a two-stage approach, initially treating the problem as a language modeling task with the prediction of the next token in SMILES strings. Afterwards, it utilizes RL to optimize chemical features, notably emphasizing qualities such as a quantitative estimate of QED (Quantitative Estimate of Drug-likeness). This demonstrates the smooth integration of GPT’s language processing abilities with the accuracy of RL approaches.
@@ 4.1.3. Encoder-Decoder Models
Dataset Time Subset/Version Scale Keywords/Tasks
@@
PCQM4M 2021.10 PCQM4Mv2 (Hu et al., 2021) 3.7M Quantum property, molecular graphs pred.
@@
Benchmarks MoleculeNet (Wu et al., 2018) 2017.10 ESOL 1k Physical chemistry prediction
FreeSoly 1K
@@
PCBA 437K Biophysics prediction
MUV 92K
HIV 41K
PDBbind 11K
BACE 1K
@@
Tox21 7K
ToxCast 8K
SIDER 1K
ClinTox 1K
@@ 4.2. Datasets
  • PCQM4M (Hu et al., 2021; Ying et al., 2021). The PCQM4Mv2 and PCQM4M-LSC datasets are designed for graph-level prediction tasks in quantum chemistry, primarily focusing on predicting the Density Functional Theory (DFT) calculated HOMO-LUMO energy gaps of molecules from their 2D molecular graphs. These datasets provide a robust platform for developing and benchmarking advanced machine learning models in predicting molecular properties. The primary difference between PCQM4Mv2 and PCQM4M-LSC is that PCQM4Mv2 is an enhanced version with more data and features, whereas PCQM4M-LSC is part of the OGB Large-Scale Challenge, emphasizing the development of state-of-the-art graph ML models with a specific scaffold data split and feature set.

### v2

@@ 3.3. Evaluation
Capabilities. The evaluation of LLMs is often structured around Bloom’s taxonomy (Krathwohl, 2002; Forehand, 2010), which outlines six cognitive levels of educational learning objectives: Remember, Understand, Apply, Analyze, Evaluate and Create. More recently, SciEval (Sun et al., 2023) has suggested assessing scientific LLMs across four dimensions: basic knowledge, knowledge application, scientific calculation, and research ability. Each dimension aligns with one or more cognitive levels in Bloom’s taxonomy. While these frameworks offer cognitive and research perspectives, our focus is to gauge the depth of knowledge grasped by LLMs, ideally comparable to human comprehension. Therefore, this survey categorizes the capabilities of Text-Sci-LLMs based on the complexity of scientific knowledge, differentiating between Pre-college, College, and Post-college levels.
@@
  • College Level. Scientific knowledge at the college level is typically more specialized, abstract, and necessitates logical reasoning and proof rather than mere memorization. It undergoes frequent updates aligned with scientific and industrial advancements, posing a challenge for LLMs to grasp and adapt to these continual changes. The master of specialized knowledge is designated as the college level of LLMs, which typically aligns with the Apply and Analyze levels in Bloom’s taxonomy, as well as the knowledge application and the scientific calculation in SciEval. Most evaluation benchmarks such as PubMedQA (Jin et al., 2019b) and SciQ (Johannes Welbl, 2017) focus on this level.
@@
  • LLM-Score involves using an LLM with a carefully designed prompt to evaluate the quality of generated text. The LLM is prompted to assess the coherence, relevance, and accuracy of the output compared to the reference answer.
When the ground-truth is absent, one could sort to experts or strong LLMs like GPT-4 (OpenAI, 2023) for assessing LLMs. Human evaluation plays a pivotal role in assessing the quality and accuracy of model-generated results, as it closely aligns with real-world application scenarios and provides comprehensive and precise feedback. However, due to extensive human involvement, this process is time-consuming and labor-intensive. Therefore, there is a growing tendency that employ GPT-4 as a “judge” to evaluate different models (Bubeck et al., 2023). Overall, these metrics are often used complementarily in various NLP tasks to provide a comprehensive evaluation of LLMs.
@@ 4. Molecular Large Language Models
(Sec. 4.1),text width=4em [Encoder-only, text width= 5.5em [BERT-based: SMILES-BERT (Wang et al., 2019)   MTL-BERT (Zhang et al., 2022c)  MolBERT (Fabian et al., 2020)
rxnfp-BERT (Schwaller et al., 2021)  Mol-BERT (Li and Jiang, 2021)  MolFormer (Ross et al., 2022)   MolRoPE-BERT (Liu et al., 2023e) ] [RoBERTa-based: ChemBERTa (Chithrananda et al., 2020)  MFBERT (Abdel-Aty and Gould, 2022)
SELFormer (Yüksel et al., 2023)   Semi-RoBERTa (Tran and Ekenna, 2023)  ] [Integrating 2D or 3D graphs: GROVER (Rong et al., 2020)  MAT (Maziarka et al., 2020)  MG-BERT (Zhang et al., 2021)  R-MAT (Maziarka et al., 2024)
KPGT (Li et al., 2022)  AGBT (Chen et al., 2021)   Molformer (Wu et al., 2023b)  Uni-Mol (Zhou et al., 2023a)  GTMGC (Xu et al., [n. d.])   ] ] [Decoder-only, text width= 5.5em [GPT-based: MolGPT (Bagal et al., 2021)  SMILESGPT (Adilov, 2021)  cMolGPT (Wang et al., 2023j)   Taiga (Mazuz et al., 2023)  iupacGPT (Cho et al., 2023) ] ] [Encoder-decoder, text width=5.5em [Vanilla Transformer-based:
@@
(Sec. 4.2),text width=4em [Pre-training, text width=4em [ZINC (Sterling and Irwin, 2015)  PubChem (Wang et al., 2012)  USPTO (Schneider et al., 2016)  PCQM4M (Hu et al., 2023; Ying et al., 2021)  GEOM (Axelrod and Gomez-Bombarelli, 2022)
MolTLU (Beaini et al., 2023)  ChEMBL (Gaulton et al., 2012)   DrugBank (Wishart et al., 2018)  GDB-17 (Ruddigkeit et al., 2012)  ExCAPE-DB (Sun et al., 2017)  ] ] [Benchmarks, text width=4em [MoleculeNet (Wu et al., 2018)  MARCEL (Zhu et al., 2023)  GuacaMol (Brown et al., 2019)
MOSES (Polykovskiy et al., 2020)  ADMETlab 2.0  (Xiong et al., 2021)  Molecule3D (Xu et al., 2021)  ] ] ] [Evaluation
(Sec. 4.3),text width=4em [Property prediction,text width=7em [LogP   Solubility  Toxicity  ADMET   Lipophilicity   Stability   etc.] ] [Interaction prediction,text width=7em [Drug-drug interactions] ] [Reaction prediction,text width=7em [Forward reaction prediction and Retrosynthetic analysis] ] [Molecule generation,text width=7em [Template-based design and De novo design] ] [Evaluation metric, text width=7em [Discriminative tasks: Accuracy   Recall   F1-score  etc. ] [Generative tasks: Validity   Unique   Novelty   Frag   Scaff   IntDivp   FCD  etc. ] ] ] [Summary
@@ 4.1. Models
As Transformer (Vaswani et al., 2017) can effectively model sequential data, numerous Transformer-based Mol-LLMs have been proposed for predicting molecular properties and structures. Janakarajan et al. (Janakarajan et al., 2023), Bran and Schwaller (Bran and Schwaller, 2023) had extensively reviewed the application of Transformer technology in molecular discovery, underscoring its significant impact on advancing the field. Here we categorize Mol-LLMs into three distinct types based on their specific architectures: encoder-only, decoder-only, and encoder-decoder models, as shown in Table 3.
@@ 4.1.1. Encoder-only Models
Encoder-only models focus on understanding and interpreting the input molecules, making them well-suited for tasks that require a deep comprehension of molecular structures and properties. A typical example is SMILES-BERT (Wang et al., 2019) that leverages the BERT architecture to interpret SMILES representations of molecules. It undergoes pre-training on an extensive dataset of unlabeled data through masked SMILES recovery tasks, enabling it to learn nuanced patterns and relationships within molecular structures. Nevertheless, the fine-tuning phase for individual downstream tasks is inefficient. MTL-BERT (Zhang et al., 2022c) mitigates this limitation by employing a multitask learning framework, which substantially increases the accuracy of predictions for multiple molecular properties. MolBERT (Fabian et al., 2020) revolutionizes molecular representation learning by integrating SMILES Equivalence and physical-chemical property prediction tasks, significantly enhancing performance in virtual screening and QSAR benchmarks. The rxnfp-BERT (Schwaller et al., 2021) model draws on the principles of BERT to process and learn from chemical reactions. It is designed to generate reaction fingerprints, which are representations of chemical reactions used for various predictive tasks. However, these models predominantly concentrate on contextual information in molecular sequences while giving less attention to molecular substructures, such as functional groups. Mol-BERT (Li and Jiang, 2021) extracts atomic-level and substructural features centered on the current atom. MolFormer (Ross et al., 2022) further enhances this approach by integrating relative position encoding and rotational position embedding, thereby facilitating the comprehension of spatial relationships between atoms in a molecule. MolRoPE-BERT (Liu et al., 2023e) augments this method by employing Rotary Position Embedding (RoPE) to efficiently encode positional information in SMILES, aiming to provide a more robust understanding of molecules.
@@
MolBERT (Fabian et al., 2020) 2020.11 85M BERT ChEMBL Prop. pred.
@@
R-MAT (Maziarka et al., 2024) 2021.10 - Trans. enc. ZINC-15, ChEMBL Prop. pred.
@@
MTL-BERT (Zhang et al., 2022c) 2022.12 - BERT ChEMBL Prop. pred.
MolRoPE-BERT (Liu et al., 2023e) 2023.01 - BERT ZINC-15, ChEMBL 27 Prop. pred. $\times$
Molformer (Wu et al., 2023b) 2023.01 - Trans. enc. ZINC, PubChem Prop. pred.
@@
cMolGPT (Wang et al., 2023j) 2023.05 - GPT MOSES Mol. gen.
@@
The above BERT-based models integrate a vast amount of unlabeled molecular data through a self-supervised learning strategy, but they are based on the sequential format of molecule SMILES or SELFIES, overlooking crucial graph structures and three-dimensional (3D) chemical information. The oversight of graph structures can be effectively addressed by integrating graphs with Transformers, such as GROVER (Rong et al., 2020), MAT (Maziarka et al., 2020), R-MAT (Maziarka et al., 2024), MG-BERT (Zhang et al., 2021), KPGT (Li et al., 2022) and GTMGC (Xu et al., [n. d.]). For example, MG-BERT (Zhang et al., 2021) combines GNN’s local message passing with BERT’s architecture, enhancing molecular feature extraction and utilizing a unique pretraining strategy involving atom masking. Certainly, GNNs alone can deal with molecule graph structures, and representative works include KCL (Fang et al., 2022), KANO (Fang et al., 2023d), GS-Meta (Zhuang et al., 2023b), and iMoLD (Zhuang et al., 2023a). Moreover, several approaches have been employed to model 3D molecules, including AGBT (Chen et al., 2021), Molformer (Wu et al., 2023b), and Uni-mol (Zhou et al., 2023a), and DeepSorption (Cui et al., 2023), showcasing the evolution and diversification of computational methods for molecular modeling. For example, DeepSorption leverages a SE(3)-equivalent Transformer to effectively model the 3D structures of molecular crystals, leading to accurate prediction of gas adsorption. Graph and 3D structure-based molecular representation learning is out of the scope of this survey and we refer interested readers to the surveys (Guo et al., 2023; Xia et al., 2023).
@@ 4.1.2. Decoder-only Models
In recent years, decoder-only architectures like GPT have revolutionized molecular generation in cheminformatics. Mol-LLMs based on the GPT architecture, including MolGPT (Bagal et al., 2021), SMILESGPT (Adilov, 2021), iupaccGPT (Cho et al., 2023), cMolGPT (Wang et al., 2023j), and Taiga (Wang et al., 2022b), primarily use SMILES strings as input, effectively addressing the challenge of navigating the vast chemical space. These models are pivotal in drug discovery and material science, enabling the synthesis of molecules with specific properties.
MolGPT (Bagal et al., 2021), a pioneer in utilizing GPT for molecular generation, uses conditional training for property optimization. It stands out for its efficiency and effectiveness in molecular modeling and drug discovery, showing strong control over multiple properties for accurate generation. Moreover, both SMILESGPT and iupacGPT are based on the GPT-2 architecture. SMILESGPT (Adilov, 2021) is a causative Transformer tailored for drug discovery, training on SMILES notation with BPE, and employing special embedding layers for effective molecular representation learning. iupacGPT (Cho et al., 2023) diverges by using the more readable IUPAC nomenclature, excelling in learning hierarchical and inner-sequence relationships between atoms and chemical groups. The cTransformer (Wang et al., 2022b) and cMolGPT (Wang et al., 2023j) models are specifically developed for the purpose of generating molecules based on given conditions. cTransformer (Wang et al., 2022b) synergizes the Transformer architecture with conditional embeddings, allowing for target-specific molecule generation and fine-tuning with target-specific data. cMolGPT (Wang et al., 2023j) represents an advancement in this domain by generating target-specific compounds by utilizing target embeddings as keys and values within its multi-head attention, adeptly producing molecules active against designated targets. Taiga (Mazuz et al., 2023) exemplifies the combination of GPT and reinforcement learning (RL) methodologies. It employs a two-stage approach, initially treating the problem as a language modeling task with the prediction of the next token in SMILES strings. Afterwards, it utilizes RL to optimize chemical features, notably emphasizing qualities such as a quantitative estimate of QED (Quantitative Estimate of Drug-likeness). This demonstrates the smooth integration of GPT’s language processing abilities with the accuracy of RL approaches.
@@ 4.1.3. Encoder-Decoder Models
Dataset Last updated Subset/Version Scale Keywords/Tasks
@@
PCQM4M 2021.10 PCQM4Mv2 (Hu et al., 2023) 3.7M Quantum property, molecular graphs pred.
@@
Benchmarks MoleculeNet (Wu et al., 2018) 2017.10 QM7 7K Quantum mechanics prediction
QM8 22K
QM9 134K
ESOL 1K Physical chemistry prediction
FreeSoly 643
@@
PCBA 440K Biophysics prediction
MUV 93K
HIV 42K
PDBbind 12K
BACE 1.5K
@@
Tox21 8K
ToxCast 8.6K
SIDER 1.4K
ClinTox 1.4K
@@
SPECTRA  (Ektefaie et al., 2024) 2024.02 - 637k Generalization ability evaluation
Molecule3D (Xu et al., 2021) 2021.09 GEOM, QM9 3.9M Ground-state 3D Molecular Geometry Prediction
@@ 4.2. Datasets
  • PCQM4M (Hu et al., 2023; Ying et al., 2021). The PCQM4Mv2 and PCQM4M-LSC datasets are designed for graph-level prediction tasks in quantum chemistry, primarily focusing on predicting the Density Functional Theory (DFT) calculated HOMO-LUMO energy gaps of molecules from their 2D molecular graphs. These datasets provide a robust platform for developing and benchmarking advanced machine learning models in predicting molecular properties. The primary difference between PCQM4Mv2 and PCQM4M-LSC is that PCQM4Mv2 is an enhanced version with more data and features, whereas PCQM4M-LSC is part of the OGB Large-Scale Challenge, emphasizing the development of state-of-the-art graph ML models with a specific scaffold data split and feature set.
@@
  • SPECTRA (Ektefaie et al., 2024). The SPECTRA benchmark is a cutting-edge framework designed to assess the generalizability of AI models across molecular and biological datasets by evaluating their performance on diverse training and testing splits. It illuminates the robustness of models in predicting unseen data, showcased through its application to a variety of datasets from tuberculosis to SARS-CoV-2. This targeted analysis helps in refining AI approaches for precise applications in molecular biology and bioinformatics, ensuring the development of more accurate and specialized AI models.
  • Molecule3D (Xu et al., 2021). Molecule3D is a benchmark dataset specifically curated for predicting ground-state 3D molecular geometries from molecular graphs, containing approximately 3.9 million molecules with precise geometries obtained via density functional theory (DFT). This dataset facilitates the development and evaluation of machine learning models in capturing the 3D spatial configurations of molecules, which are crucial for accurate molecular property prediction and other downstream tasks.